The amount of money that a salesperson makes varies directly as the total amount of sales made. If the salesperson $7,500 for selling a house for $120,000, how much will he or she make if he or she sells a house for $150,000?
Use direct variation to solve problem

Answer by tutorcecilia(2152)

$7,500 commission/$120,000 cost of house = x commission/$150,000 cost of house
7,500/120,000=x/150,000 [cross-multiply]
(7500)(15,000)=120,000x [solve for x]
1,125,000,000/120,000=x
x=$9,375 commission